a cube of ice floats partly in water and partly in oil .find the ratio of the volume of ice immersed in water to thatin oil. specific gravity of oil is 0.8 and that of ice is 0.9

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear aastha

wieght of ice cube should be equal to the sum of wieght of oil replaced and  wight of water replaced by ice cube

let the volume in water is v1 and in oil is v2

0.9 * g * (v1+v2) =1* g*v1 + 0.8 * g* v2

or  0.9(v1/v2 +1) = v1/v2 +0.8

or v1/v2 =1


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin